OpenStudy (anonymous):

Which would be bigger , a 26 inch pizza or two 13 inch pizzas ? Or would they be the same?

OpenStudy (anonymous):

they are both the same

OpenStudy (anonymous):

so it would be equal

OpenStudy (precal):

use the area of a circle \[A=\pi*r^2\]

OpenStudy (mertsj):

A 26 inch pizza is bigger.

OpenStudy (anonymous):

^^they are both the same

OpenStudy (anonymous):

They are not the same.

OpenStudy (anonymous):

\[\frac{13^2\pi}{2*6.5^2\pi}\]

OpenStudy (anonymous):

if u multiply 13 and 2 together u get 26

OpenStudy (anonymous):

Yes, but that's now how you solve it. You need to use a ratio of areas of circles.

OpenStudy (anonymous):

o so A=π∗r2

OpenStudy (anonymous):

The 26in pizza is actually equivalent to 4 13 in pizzas.